# Convergence analysis of Levenberg-Marquardt method with Singular Scaling for non-zero residue nonlinear least-squares problems [^1] This work is supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]

### v1

@@ 7 Final remarks
We have investigated convergence issues of Levenberg-Marquardt method with singular scaling (LMMSS) [5] for nonlinear least-squares problems with non-zero residue. Through a couple of modifications in the algorithm proposed in [5, Algorithm 1], mainly, the choice of Levenberg-Marquardt parameter and the line-search, we have established local and global convergence results under mild assumptions.
It was shown that LMMSS direction is well defined as long as the nullspace of the Jacobian and the nullspace of the singular scaling matrix $L^{T}L$ have trivial intersection (Assumption 1). Then, it was demonstrated local superlinear convergence rate under a local error bound condition on the gradient of the least-squares function (Assumption 3) and an assumption on the linearization error of the gradient (Assumption 4) which represents a measure of nonlinearity and residue size. These results were established for both the constant rank and diminishing rank cases.
By using an Armijo line-search and safeguard directions, it was proved that all limit points of the sequence generated by Algorithm 1 are stationary points of the least-squares function. We also presented some examples illustrating the assumptions and convergence results.
In future works we plan to consider Hölder metric subregularity [1] instead of the local error bound used here and address local convergence when a weaker form of Assumption 4 holds, namely when $\|J(x)^{T}F(z)\|={\cal O}(\|x-z\|)$ instead of $\|J(x)^{T}F(z)\|=o(\|x-z\|)$, following the lines of the analysis presented in [2].
Apart from tackling these questions, we aim to conduct computational experiments to test the algorithm, equipped with suitable stopping criteria, in certain inverse problems to assess its numerical behavior.
@@ Acknowledgments
This work was supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]. The work of DG was partially supported by CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico), Brazil [grant number 305213/2021-0].

### v2

@@ 7 Final remarks
We have investigated the local convergence of Levenberg-Marquardt method with singular scaling (LMMSS) proposed in Boos2024 when applied to nonlinear least-squares problems with nonzero residual.
Our study reveals that, regardless the Jacobian rank is constant or not in a neighborhood of $x^{*}\in X^{*}$, the convergence rate of $\textup{dist}(x_{k},X^{*})$ to zero depends on a combined measure of nonlinearity and residual size $\|J(x)^{T}F(z)\|$ (see Assumption 3 and 4) and suitable choices of the LM parameter $\lambda_{k}$ (Theorems  25). For $\|J(x)^{T}F(z)\|=o(\|x-z\|)$, the local convergence rate is superlinear whereas $\|J(x)^{T}F(z)\|\leq\sigma\|x-z\|$ leads to linear convergence, but only when $\sigma$ is small enough relatively to the error bound constant $\omega$. These results are aligned with those of behling2019local but consider the more general case where $L^{T}L$ can be singular, as long as the uniform completeness condition (Definition 1) holds in a suitable neighborhood of $x^{*}\in X^{*}$.
Furthermore, through a few modifications in the algorithm proposed in (Boos2024, , Algorithm 1), namely, the choice of LM parameter, the full step acceptance criterion and the use of classic LMM direction as a safeguard whenever LMMSS fails to satisfy descent properties; we have established under mild assumptions the global convergence of Algorithm 1: every limit point of the generated sequence is stationary for the nonzero residual nonlinear least squares problem.
In future works we will look for weaker assumptions to control $\|J(x)^{T}F(z)\|$. The condition $\|J(x)^{T}F(z)\|\leq\sigma\|x-z\|$ with small enough $\sigma$ may be too restrictive and hold only in nonzero residual problems with very small residual norm. When it does not hold, we cannot even ensure that the “pure” LMMSS converges locally. In that case, other methods such as Quasi-Newton or even Newton, when the Hessians of $F_{i}(x)$ are available, might be more appropriate.
@@ Acknowledgments
This work was supported by Brazilian agencies FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360] and CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico) [grant number 305213/2021-0].